Disposal of fixed asset

Instructions

Chart of Accounts

First Questions

Journal

X

Instructions

Equipment acquired on January 6 at a cost of $417,390, has an estimated useful life of 17 years and an estimated residual value of $68,805.

A. What was the annual amount of depreciation for the Years 1-3 using the straight-line method of depreciation?
B. What was the book value of the equipment on January 1 of Year 4?
C. Assuming that the equipment was sold on January 3 of Year 4 for $339,855, journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.
D.

Assuming that the equipment had been sold on January 3 of Year 4 for $368,885 instead of $339,855, journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.

CHART OF ACCOUNTS
General Ledger
ASSETS
110 Cash
111 Petty Cash
112 Accounts Receivable
114 Interest Receivable
115 Notes Receivable
116 Merchandise Inventory
117 Supplies
119 Prepaid Insurance
120 Land
123 Delivery Truck
124 Accumulated Depreciation-Delivery Truck
125 Equipment
126 Accumulated Depreciation-Equipment
130 Mineral Rights
131 Accumulated Depletion
132 Goodwill
133 Patents
LIABILITIES
210 Accounts Payable
211 Salaries Payable
213 Sales Tax Payable
214 Interest Payable
215 Notes Payable
EQUITY
310 Common Stock
311 Retained Earnings
312 Dividends
313 Income Summary
REVENUE
410 Sales
610 Interest Revenue
620 Gain on Sale of Delivery Truck
621 Gain on Sale of Equipment
EXPENSES
510 Cost of Merchandise Sold
520 Salaries Expense
521 Advertising Expense
522 Depreciation Expense-Delivery Truck
523 Delivery Expense
524 Repairs and Maintenance Expense
529 Selling Expenses
531 Rent Expense
532 Depreciation Expense-Equipment
533 Depletion Expense
534 Amortization Expense-Patents
535 Insurance Expense
536 Supplies Expense
539 Miscellaneous Expense
710 Interest Expense
720 Loss on Sale of Delivery Truck
721 Loss on Sale of Equipment

A. What was the annual amount of depreciation for the Years 1-3 using the straight-line method of depreciation?

Year 1 depreciation expense
Year 2 depreciation expense
Year 3 depreciation expense

B. What was the book value of the equipment on January 1 of Year 4?

C. Assuming that the equipment was sold on January 3 of Year 4 for $339,855, journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 1

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

2

3

4

D. Assuming that the equipment had been sold on January 3 of Year 4 for $368,885 instead of $339,855, journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 1

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

2

3

4
Answer #1

Answer:-

A. Depreciation per year = (Total cost - Residual value) ÷ Useful life = (417,390-68,805)÷17 = $20,505

2013 Depreciation expense $20,505
2014 Depreciation expense $20,505
2015 Depreciation expense $20,505

B. Total depreciation for three years = 20,505×3 = $61,515

Book value of the equipment on January 1, 2016 = Total cost - Total depreciation for three years = 417,390-61,515 = $355,875

C.

Journal
Date Description Post. Ref. Debit Credit
Jan. 3 Cash 339,855
Loss on equipment disposal (bal. fig.) 16,020
Equipment (book value) 355,875
(To record sale of equipment at $339,855, which is lower than its book value, incurring a loss.)

D.

Journal
Date Description Post. Ref. Debit Credit
Jan. 3 Cash 368,885
Gain on equipment disposal (bal. fig.) 13,010
Equipment (book value) 355,875
(To record sale of equipment at $368,885, which is higher than its book value, incurring a gain.)

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

                                                *******THANK YOU*******